A capacitor with stored energy 4.0 J is connected with an identical capacitor with no electric field in between. Find the total energy stored in the two capacitors.

Grade:11

1 Answers

Jitender Pal
9 years ago
Sol. Stored energy of capacitor C base 1 = 4.0 J = 1/2 q^2/c^2 = 4.0 J When then connected, the charge shared 1/2 q base 1^2/c^2 = 1/2 q base 2^2/c^2 ⇒ q base 1 = q base 2 So that the energy should divided. ∴ The total energy stored in the two capacitors each is 2 J.